OpenStudy (anonymous):

simplify 4-8x3/4+1/4 @Mertsj @e.cociuba @chrismoon

OpenStudy (anonymous):

thts not a choice

OpenStudy (anonymous):

a.-7/4 b.8 c.11/4 d.3

OpenStudy (anonymous):

i need 4 more anwsered in 20 min after this can u help me fast please

OpenStudy (mertsj):

\[4-8\times \frac{3}{4}+\frac{1}{4}=4-6+\frac{1}{4}=-2+\frac{1}{4}=-1\frac{3}{4}\]

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

dan has 3.75 if he spends 1/3 of his money how much will he have left

OpenStudy (mertsj):

If sh spends 1/3 he has 2/3 left so multiply 2/3 by 3.75

OpenStudy (mertsj):

2.50 is what you got?

OpenStudy (mertsj):

What are the choices?